OpenStudy (anonymous):

Find the exact value of the expression. a) cos(arcsin(8/17)) b) sin^-1(tan(pi/4))

OpenStudy (anonymous):



OpenStudy (anonymous):

there is the picture of an angle whose sine is \(\frac{8}{17}\) to find the cosine of that angle, you need the other side of the triangle by pythagoras you get \[a=\sqrt{17^2-8^2}=15\]

OpenStudy (anonymous):



OpenStudy (anonymous):

now you can see that \(\cos(\theta))=\cos(\sin^{-1}(\frac{8}{17}))=\frac{15}{17}\)